# The Active and Noise-Tolerant Strategic Perceptron

### v1

@@ 2 Model and Preliminaries
We denote the instance space by $\cal Z$ and the label space by $\cal Y$. The true feature vectors $\bm{z}_{t}$ belong to instance space ${\cal Z}=\{\bm{z}\in R^{d}:\|\bm{z}\|\leq 1\}$; a unit $d$-dimensional ball. The label space ${\cal Y}=\{+1,-1\}$. We assume all examples $\bm{z}$ are drawn i.i.d. from the uniform distribution $D$ over $\cal Z$. Upon sampling an example, our algorithm observes $\bm{x}$, whose true instance vector $\bm{z}\in\cal Z$ is drawn from $D$ and whose label is hidden by default. Our algorithm is allowed to make queries to a labeling oracle $\cal O$, which returns the true label for $\bm{z}$. In line with prior work (Dasgupta2005Analysis; DBLP:conf/nips/YanZ17) on nonstrategic settings, the goal of the learning algorithm is to classify the true instance vectors accurately while minimizing the number of label queries. To achieve this, we leverage active learning techniques, which allow querying labels only when necessary, reducing reliance on labeled data from a costly oracle.
@@ 3 Active and Noise-Tolerant Strategic Perceptron
3.1 The Non-Strategic Active Perceptron of DBLP:conf/nips/YanZ17
Our starting point is the algorithm proposed by Yan and Zhang DBLP:conf/nips/YanZ17. We begin by outlining its main ideas before introducing our modifications for the strategic setting. The algorithm is organized into an outer and an inner layer, similar in structure to Algorithm 1 and Algorithm 2, and proceeds in a sequence of epochs. The outer layer initializes with a hypothesis vector $\bm{v}_{0}$ and invokes the inner layer in successive epochs, each with updated parameters such as the target error, confidence level, and active learning bandwidth. The outcome of each epoch is an updated hypothesis $\bm{v}_{i}$, which serves as the starting point for the next. The total number of epochs is logarithmic in $\nicefrac{{1}}{{\varepsilon}}$, where $\varepsilon$ is the final target error.
The inner layer defines both the update rule and the label-query mechanism. In the non-strategic setting, it specifies a label-query region $R_{t}$; in the implementation of DBLP:conf/nips/YanZ17, which assumes that examples lie on the unit sphere, this region is the set of points whose dot product with the current hypothesis lies in $\nicefrac{{b}}{{2}},b]$. Compared to earlier active Perceptron algorithms such as Dasgupta2005Analysis, the use of a lower bound on the dot product ensures that each update makes sufficient progress, thereby accelerating convergence. As for the update rule, when the algorithm makes a mistake on a queried example, it updates the current hypothesis using $\bm{w}_{t+1}=\bm{w}_{t}\pm 2(\bm{w}_{t}\cdot\bm{x}_{t})\bm{x}_{t}$. This update rule, first proposed by Dasgupta2005Analysis, guarantees that the angle between the current hypothesis and the optimal separator $\bm{u}$ monotonically decreases. Furthermore, it preserves the unit norm of the hypothesis vector as long as $|\bm{x}_{t}|=1$.
@@ 3.2 Our Strategic Variant of the Active Perceptron
Similar to DBLP:conf/nips/YanZ17, we consider an outer layer, Algorithm 1, and inner layer, Algorithm 2, for our algorithm, where the new prediction rule, label query region, and update rule are incorporated in the inner layer.
@@
Rather than using the standard threshold $\bm{v}_{t}\cdot\bm{x}_{t}\geq 0$, our prediction rule raises the threshold to $\bm{v}_{t}\cdot\bm{x}_{t}\geq\frac{1}{c}$, where $c$ is the cost per unit of manipulation. This adjustment, following DBLP:conf/sigecom/AhmadiBBN21, accounts for agents’ strategic behavior and ensures that, conditioned on convergence to the optimal classifier, (the majority of the) truly positive points either lie on the positive side or can manipulate to reach it. Meanwhile, truly negative points remain on the negative side and would incur negative utility if they attempted to manipulate and be classified as positive.
@@ 3.3 Proof of the Main Result
Next, we invoke a result from DBLP:conf/nips/YanZ17 that shows in the nonstrategic setting, the difference in the error rate of two prediction rules can be related to the angles of their normal vectors.
Lemma 5 (DBLP:conf/nips/YanZ17 Lemma 1).
@@
Theorem 9 (Adapted from (DBLP:conf/nips/YanZ17), Theorem 3).
@@
Lemma 10 (Adapted from (DBLP:conf/nips/YanZ17), Lemma 3).
@@
Similar to the proof of Theorem 3 in (DBLP:conf/nips/YanZ17), we define, for each iteration $k$, a corresponding event with high individual success probability. Specifically, by Lemma 10, for every $k$, there exists an event $E_{k}$ such that $\Pr(E_{k})\geq 1-\frac{\delta}{k(k+1)}$. Moreover, on event $E_{k}$, items 1 through 4 of Lemma 10 hold for the input $\bm{w}_{0}=\bm{v}_{k}$.
@@ 4 Discussion
Much of the prior literature on active classification, e.g., Dasgupta2005Analysis; DBLP:conf/nips/YanZ17, assumed for simplicity and cleaner mathematical formulations, that all examples lie on the surface of a unit ball; for illustration, the circle in Figure 1. While this assumption was inconsequential to the goals of previous research, it becomes crucial in strategic scenarios, where the relationship between observed and true features is strongly influenced by the geometry of the space. In the strategic setting, this assumption implies a one-to-one mapping between observed feature vectors, $\bm{x}_{i}$ in Figure 1 and original feature vectors, $\bm{z}_{i}$, and provides a straightforward case for analysis, as it allows the original unmanipulated positions of examples to be completely recovered. Specifically, given a linear classifier, e.g., $\bm{v}\cdot\bm{x}\geq 1/c$, and observing an example at position $\bm{x}$, the true position $\bm{z}$ of the example on the surface of the unit ball can be recovered through an orthogonal projection, leveraging properties of the utility function, as shown in Lemma 3.
@@
Figure 1: Illustration of observed regions of examples, where the original examples are on the surface of a unit ball, as assumed in prior literature on active learning Dasgupta2005Analysis; DBLP:conf/nips/YanZ17. The colored area shows the potential positions where the examples can be observed, following Lemma 3. Under this assumption, there is a one-to-one mapping between the original examples and observed ones, and observing an example $\bm{x}_{i}$ exactly identifies the original position $\bm{z}_{i}$. For instance, when observing $\bm{x}_{1}$ to recover the original point $\bm{z}_{1}$, one needs to consider the projection of $\bm{x}_{1}$ on the surface of the pall, in the opposite direction of $\bm{v}$. This one-to-one mapping no longer exists under the relaxed assumptions that original examples can be inside the ball and not just the surface.
@@
Although we have focused on the active learning algorithm by DBLP:conf/nips/YanZ17 as our baseline, and presented a modified version of this algorithm for the strategic setting, our ideas have the potential to the be applied to other such algorithms. Particularly, following the new prediction rule and label query region, tailored to strategic setting, our analysis provides a reduction from the strategic scenario to non-strategic ones that could be applicable to a variety of active classification algorithms.
@@
Finally, we conjecture using similar ideas as DBLP:conf/sigecom/AhmadiBBN21, our algorithms can be generalized to scenarios where the learner does not know the manipulation costs of strategic agents upfront and needs to learn throughout the algorithm, and scenarios with different utility and cost functions, e.g., weighted $\ell_{1}$ costs of manipulation.
@@ Appendix A Missing Proofs Related to Strategic Behavior
As implied from Observation 1 in [DBLP:conf/sigecom/AhmadiBBN21], the agent with attributes vector $\bm{z}_{t}$ only manipulates iff by movement cost at most $c$ can be classified as positive. We now analyze the three cases: (i) If $\bm{z}_{t}\cdot\bm{v}_{t}<0$, any movement incurs negative utility, so the agent does not move. (ii) If $0\leq\bm{z}_{t}\cdot\bm{v}_{t}<1/c$, the agent maximizes utility by moving to $\bm{x}\cdot\bm{v}_{t}=1/c$ in direction $\bm{v}_{t}$, i.e.,
@@ Appendix B Initialization Step
Algorithm 1 assumes that the initial vector $\bm{v}_{0}$ forms an angle less than $\pi/2$ with the true separator $\bm{u}$. To obtain such a vector, we adopt an approach similar to that of [DBLP:conf/nips/YanZ17, DBLP:conf/stoc/AwasthiBL14]. Intuitively, Algorithm 3 begins with two opposite vectors and, after a small number of trials, selects the one with lower classification error. This initialization step incurs only a constant overhead in terms of label, time, and mistake complexities.
@@ 4    $S=S\cup\{\bm{x}\}$
Algorithm 3 Master Algorithm in Adversarial Noise Setting, (adapted from [DBLP:conf/nips/YanZ17, DBLP:conf/stoc/AwasthiBL14])
By Corollary 7 and Theorem 2, we can effectively disregard the strategic behavior of agents and apply Theorem 12 from [DBLP:conf/nips/YanZ17], which implies the following.

### v2

@@ 2 Model and Preliminaries
We denote the instance space by $\cal Z$ and the label space by $\cal Y$. The true feature vectors $\bm{z}_{t}$ belong to instance space ${\cal Z}=\{\bm{z}\in R^{d}:\|\bm{z}\|\leq 1\}$; a unit $d$-dimensional ball. The label space ${\cal Y}=\{+1,-1\}$. We assume all examples $\bm{z}$ are drawn i.i.d. from the uniform distribution $D$ over $\cal Z$. Upon sampling an example, our algorithm observes $\bm{x}$, whose true instance vector $\bm{z}\in\cal Z$ is drawn from $D$ and whose label is hidden by default. Our algorithm is allowed to make queries to a labeling oracle $\cal O$, which returns the true label for $\bm{z}$. In line with prior work (DKM, 05; YZ, 17) on nonstrategic settings, the goal of the learning algorithm is to classify the true instance vectors accurately while minimizing the number of label queries. To achieve this, we leverage active learning techniques, which allow querying labels only when necessary, reducing reliance on labeled data from a costly oracle.
@@ 3 Active and Noise-Tolerant Strategic Perceptron
3.1 The Non-Strategic Active Perceptron of YZ (17)
Our starting point is the algorithm proposed by Yan and Zhang YZ (17). We begin by outlining its main ideas before introducing our modifications for the strategic setting. The algorithm is organized into an outer and an inner layer, similar in structure to Algorithm 1 and Algorithm 2, and proceeds in a sequence of epochs. The outer layer initializes with a hypothesis vector $\bm{v}_{0}$ and invokes the inner layer in successive epochs, each with updated parameters such as the target error, confidence level, and active learning bandwidth. The outcome of each epoch is an updated hypothesis $\bm{v}_{i}$, which serves as the starting point for the next. The total number of epochs is logarithmic in $\nicefrac{{1}}{{\varepsilon}}$, where $\varepsilon$ is the final target error.
The inner layer defines both the update rule and the label-query mechanism. In the non-strategic setting, it specifies a label-query region $R_{t}$; in the implementation of YZ (17), which assumes that examples lie on the unit sphere, this region is the set of points whose dot product with the current hypothesis lies in $\nicefrac{{b}}{{2}},b]$. Compared to earlier active Perceptron algorithms such as DKM (05), the use of a lower bound on the dot product ensures that each update makes sufficient progress, thereby accelerating convergence. As for the update rule, when the algorithm makes a mistake on a queried example, it updates the current hypothesis using $\bm{w}_{t+1}=\bm{w}_{t}\pm 2(\bm{w}_{t}\cdot\bm{x}_{t})\bm{x}_{t}$. This update rule, first proposed by DKM (05), guarantees that the angle between the current hypothesis and the optimal separator $\bm{u}$ monotonically decreases. Furthermore, it preserves the unit norm of the hypothesis vector as long as $|\bm{x}_{t}|=1$.
@@ 3.2 Our Strategic Variant of the Active Perceptron
Similar to YZ (17), we consider an outer layer, Algorithm 1, and inner layer, Algorithm 2, for our algorithm, where the new prediction rule, label query region, and update rule are incorporated in the inner layer.
@@
Rather than using the standard threshold $\bm{v}_{t}\cdot\bm{x}_{t}\geq 0$, our prediction rule raises the threshold to $\bm{v}_{t}\cdot\bm{x}_{t}\geq\frac{1}{c}$, where $c$ is the cost per unit of manipulation. This adjustment, following ABBN (21), accounts for agents’ strategic behavior and ensures that, conditioned on convergence to the optimal classifier, (the majority of the) truly positive points either lie on the positive side or can manipulate to reach it. Meanwhile, truly negative points remain on the negative side and would incur negative utility if they attempted to manipulate and be classified as positive.
@@ 3.3 Proof of the Main Result
Next, we invoke a result from YZ (17) that shows in the nonstrategic setting, the difference in the error rate of two prediction rules can be related to the angles of their normal vectors.
Lemma 5 (YZ (17) Lemma 1).
@@
Theorem 9 (Adapted from (YZ, 17), Theorem 3).
@@
Lemma 10 (Adapted from (YZ, 17), Lemma 3).
@@
Similar to the proof of Theorem 3 in (YZ, 17), we define, for each iteration $k$, a corresponding event with high individual success probability. Specifically, by Lemma 10, for every $k$, there exists an event $E_{k}$ such that $\Pr(E_{k})\geq 1-\frac{\delta}{k(k+1)}$. Moreover, on event $E_{k}$, items 1 through 4 of Lemma 10 hold for the input $\bm{w}_{0}=\bm{v}_{k}$.
@@ 4 Discussion
Much of the prior literature on active classification, e.g., DKM (05); YZ (17), assumed for simplicity and cleaner mathematical formulations, that all examples lie on the surface of a unit ball; for illustration, the circle in Figure 1. While this assumption was inconsequential to the goals of previous research, it becomes crucial in strategic scenarios, where the relationship between observed and true features is strongly influenced by the geometry of the space. In the strategic setting, this assumption implies a one-to-one mapping between observed feature vectors, $\bm{x}_{i}$ in Figure 1 and original feature vectors, $\bm{z}_{i}$, and provides a straightforward case for analysis, as it allows the original unmanipulated positions of examples to be completely recovered. Specifically, given a linear classifier, e.g., $\bm{v}\cdot\bm{x}\geq 1/c$, and observing an example at position $\bm{x}$, the true position $\bm{z}$ of the example on the surface of the unit ball can be recovered through an orthogonal projection, leveraging properties of the utility function, as shown in Lemma 3.
@@
Figure 1: Illustration of observed regions of examples, where the original examples are on the surface of a unit ball, as assumed in prior literature on active learning DKM (05); YZ (17). The colored area shows the potential positions where the examples can be observed, following Lemma 3. Under this assumption, there is a one-to-one mapping between the original examples and observed ones, and observing an example $\bm{x}_{i}$ exactly identifies the original position $\bm{z}_{i}$. For instance, when observing $\bm{x}_{1}$ to recover the original point $\bm{z}_{1}$, one needs to consider the projection of $\bm{x}_{1}$ on the surface of the pall, in the opposite direction of $\bm{v}$. This one-to-one mapping no longer exists under the relaxed assumptions that original examples can be inside the ball and not just the surface.
@@
Although we have focused on the active learning algorithm by YZ (17) as our baseline, and presented a modified version of this algorithm for the strategic setting, our ideas have the potential to the be applied to other such algorithms. Particularly, following the new prediction rule and label query region, tailored to strategic setting, our analysis provides a reduction from the strategic scenario to non-strategic ones that could be applicable to a variety of active classification algorithms.
@@
Finally, we conjecture using similar ideas as ABBN (21), our algorithms can be generalized to scenarios where the learner does not know the manipulation costs of strategic agents upfront and needs to learn throughout the algorithm, and scenarios with different utility and cost functions, e.g., weighted $\ell_{1}$ costs of manipulation.
@@ Appendix A Missing Proofs Related to Strategic Behavior
As implied from Observation 1 in [1], the agent with attributes vector $\bm{z}_{t}$ only manipulates iff by movement cost at most $c$ can be classified as positive. We now analyze the three cases: (i) If $\bm{z}_{t}\cdot\bm{v}_{t}<0$, any movement incurs negative utility, so the agent does not move. (ii) If $0\leq\bm{z}_{t}\cdot\bm{v}_{t}<1/c$, the agent maximizes utility by moving to $\bm{x}\cdot\bm{v}_{t}=1/c$ in direction $\bm{v}_{t}$, i.e.,
@@ Appendix B Initialization Step
Algorithm 1 assumes that the initial vector $\bm{v}_{0}$ forms an angle less than $\pi/2$ with the true separator $\bm{u}$. To obtain such a vector, we adopt an approach similar to that of [28, 4]. Intuitively, Algorithm 3 begins with two opposite vectors and, after a small number of trials, selects the one with lower classification error. This initialization step incurs only a constant overhead in terms of label, time, and mistake complexities.
@@ 4    $S=S\cup\{\bm{x}\}$
Algorithm 3 Master Algorithm in Adversarial Noise Setting, (adapted from [28, 4])
By Corollary 7 and Theorem 2, we can effectively disregard the strategic behavior of agents and apply Theorem 12 from [28], which implies the following.